# DiffArtist: Towards Structure and Appearance Controllable Image Stylization

### v3

@@ 5.3. Ablations
The delegations enable dual controllability. The control mechanism of DiffArtist is built upon the two delegations. Removing either one will damage the dual controllability. Therefore we compare removing the delegations from full method with the default parameters and summarize the result in Tab. 5. The full methods achieves the best results, demonstrating the synergistic effect of delegations.
Table 5. Ablation on delegation branches. Content delegation reduces unrelated content modification, while the style delegation boosts stylization strength. Controls from both delegations are disentangled from each other.
@@
S2A injection promotes uniform spatial distribution of style strength in the style delegation process, thereby avoiding artifacts in the final stylization result. We visualize the denoised style image (from appearance delegation) and the final stylization result in Fig. 6. As observed in the examples, without the S2A injection, the appearance delegation may generate irrelevant structures with uneven appearance distribution, adversely affecting the corresponding regions of the stylization results.
Figure 6. Ablation on S2A injection. S2A injection propagate the high-level semantic to the appearance generation. It avoids spatial-misalignment of appearance-style strength.
@@ 5.4. Limitations & Future work
Disentangling the the interplay of structure and appearance in artistic styles is a long-standing challenge. Despite the progress we made in this paper, significant work remains be done. For instance, the structure control in DiffArtist is at a global level, and it cannot control the structure for each object separately. Many art styles exhibit mixed structure variation, like Surrealism and Collage art. We believe that the development of dense structure evaluators with a 2D reward signal is a promising direction (Liang et al., 2024), which would require a deeper understanding of both art and computation.
@@ 6. Conclusion
We present the first exploration of structure- and appearance-controllable image stylization. Our contributions include DiffArtist, a generator that fully disentangles structure and appearance during the diffusion process, and a human-aligned evaluator to assess structural and appearance fidelity at the semantic level. Our extensive analysis prove that semantically rich representations are essential for both structure and appearance, allowing for high-fidelity and editability during control, similar to that of a human artist. We hope these design principles broadly inspire future art generation methods to produce truly artistically meaningful paintings.
@@ Appendix A Additional Qualitative Results
Figure 7. Tradeoff between Structure and Appearance Style Control. We present the experimental (solid lines) and simulation (dotted lines) trends of structural and appearance variation in diffusion process. Due to quadratic growth, highly noisy steps are required to achieve strong appearance styles, which are associated with significant structural variation, which can violate semantics. Top: Example stylization results starting from different steps, using the prompt: “Fauvism painting”. When the appearance strength is high ($t=0.8T$), the structure (legs of the horse) are incorrectly modified.
@@
Visualization Additional appearance stylization is in Fig. 10, their source image is in Fig. 11. A grid of different image with different style is in Fig. 12. Additional structure control is in Fig. 13 and Fig. 14.
Additional qualitative comparisons on stylization. We show extended comparison with previous reference-based method in Fig. 15. More qualitative comparisons with existing text-driven image stylization and manipulation methods could be found in Fig. 16 and Fig. 17.
Additional comparisons on fine-grained control. We provide additional comparisons with other control methods in Fig. 18 and Fig. 19. These results clearly demonstrate the advantage of DiffArtist in providing disentangled structural and appearance-level style control. In particular, the Ctrl-X, as an image editing method, produce less visually pleasing result when applied to image stylization. This is because they have a different definition of appearance and structure for the editing of real photos.
@@ B.1. Theoretical analysis
(7) $\mathbf{x}_{t-1}$ $=\mathcal{A}_{t}\mathbf{x}_{t}+\mathcal{B}_{t}\epsilon_{\theta}(\mathbf{x}_{t},t;y),\quad\mathcal{A}_{t}:=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}$
$\mathcal{B}_{t}$ $:=\sqrt{1-\alpha_{t-1}}-\frac{\sqrt{\alpha_{t-1}(1-\alpha_{t-1})}}{\sqrt{\alpha_{t}}}$
@@
where $\epsilon^{\prime}$ denotes an ideal denoiser that perfectly models the transition distribution $q(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$. In Eq. 8, the conceptual denoising term from $\mathbf{x}_{T}$ to $\mathbf{x}_{\tau}$ preserves the structure and appearance in $I_{c}$. The stylization trajectory from $\mathbf{x}_{\tau}$ to $\mathbf{x}_{t}$ is what actually computed, which introduces the desired appearance based on the prompt $y$ but may also lead to uncontrolled structure generation. In this paradigm, the preservation of original appearance and generation of new structure are side-effects that are undesirable. Intuitively, larger $\tau$ leads to a shorter trajectory of $\mathbf{x}_{T}\to\mathbf{x}_{\tau}$, resulting in a stronger appearance in $\hat{\mathbf{x}}_{0}$ with weaker structure preservation (uncontrolled structure stylization), while lower $\tau$ sacrifices stylization strength for stronger structure preservation. In other words, one can not arbitrarily control the strength of appearance and structure without affecting the other.
By combining Eqn. 2 and Eqn. 8, we can quantitatively assess the strength of content preservation and stylization in diffusion process under a particular noise schedule $\alpha_{1:T}$. Specifically, we further assume the SNR of structure and appearance has linear relationship with that of the $\mathbf{x}_{t}$ for $t$:
@@ B.2. Simulation
To derive the theoretical trends of structure and appearance strength during the diffusion process, we introduce an additional assumption. Specifically, we assume that the relative significance of each unweighted denoising step $\epsilon(\mathbf{x}_{t},t;y)$ on the final stylized image remains consistent across different timesteps for both structure and appearance. In other words, we assume the SNR of structure and appearance has linear relationship with that of the $\mathbf{x}_{t}$ at time $t$ as characterized by the noise schedule $\alpha_{1:T}$:
@@
With above assumption, the effect of varying $\tau$ on structure and appearance of final stylized image could be derived in close-form. In practice, we use the following code to calculate iteratively:
@@ B.3. Empirical Result
Due to the inherent inaccuracy of DDIM inversion, the estimation of $\mathbf{x}_{\tau}$ may be imperfect, resulting in unintended modifications in the final sampled image even if no style prompt $y$ is used. To address this issue, we adopt an alternative strategy by randomly sampling 500 Gaussian noise as the $\mathbf{x}_{T}$ of content, which are paired with 500 content prompts. We treat the images that denoised using content prompts for $\tau=T$ steps as the content image, which simulate a perfect inversion technique. To stylize an image, we first denoise the $\mathbf{x}_{T}$ with content prompt for $\tau$ steps, which is subsequently denoised with the style prompt for $T-\tau$ steps. The LPIPS between the stylized image and content image is used as the empirical structural strength, whereas CLIP Deception score (correct classification rate among a set of styles) is used as the empirical appearance strength. The following 10 style prompts is used:
@@
The result for both simulation and empirical results are in Fig. 7. The result show a good fit, and it turns out that the structure modification appears to be linear, with the stylization strength is quadratic with respect to $\tau$. Moreover, this result further evidenced the issue of S-A entanglement in the diffusion process.
@@ C.1. Implementation Details
The stylized images, style prompt, and the instruction prompts are feed to MLLM for inference. We compose stylized images as a grid image with numbers at the top-left corner. The full prompt template for structure and appearance score is available in Tab. 6.
@@ C.2. Correlation with Human Preference
Human Question Collection We distributed the questionnaire on a crowd-sourcing platform, where each participant was required to complete up to 20 randomly sampled ranking tasks. An example of the user interface is provided in Fig. 20. A total of 200 participants took part in this study. To ensure the validity of the responses, we included attention-check questions. If a participant answers an attention-check question incorrectly, all of their responses will be marked invalid. Response that are made with less than 20 seconds are also removed.
@@ D.2. Additional Ablations on S2A Injection
In this section we provide additional ablations to study the effect of proposed S2A design.
@@ D.3. Failure Case
In our experiment, we identify a rare ($<$1%) and special failure case in the proposed methods. Specifically, for certain content image, its stylization result will consistently contains black and white chessboard-pattern artifacts. We provide one example in Fig. 9.

### v4

@@ 5.3. Ablations
The delegations enable dual controllability. DiffArtist’s controllability stems from delegating structure and appearance generation to separate processes. To test the necessity of each, we created two ablated variants for comparison where the structure or appearance delegation is removed. Tab. 5 presents the result of this ablation. The full methods achieves the best results, demonstrating the synergistic effect of delegations for dual controllability.
Table 5. Ablation on delegation branches. The proposed two delegations are complementary to each other, and the full methods achieves the highest fidelity.
@@
S2A injection promotes semantic-aligned spatial distribution of style strength in the style delegation process, thereby avoiding artifacts in the final stylization result. We visualize the denoised style image (from appearance delegation) and the final stylization result in Fig. 6. Without S2A injection, the appearance delegation fails to align with content semantics, generating an appearance reference image with undesired patterns and an uneven texture distribution. These flaws manifest directly in the final output as distracting visual artifacts. In contrast, the full model leverages S2A to produce a coherent style representation, resulting in a clean and high-quality final image.
Figure 6. Ablation on S2A injection. S2A injection propagates the high-level semantic to the appearance generation. It avoids spatial misalignment of appearance-style strength.
@@ 5.4. Limitations & Future work
While DiffArtist marks a significant step towards disentangled structure and appearance control, we identify several limitations that open exciting avenues for future research. For instance, the structure control in DiffArtist is at a global level, and it cannot control the structure for each object separately. Many art styles exhibit mixed structure variation, like Surrealism and Collage art. Developing dense structure evaluators with 2D feedback signals is a promising direction (Liang et al., 2024), which may be further utilized as a reward model for reinforcement learning (Black et al., 2023; Clark et al., 2023).
@@ 6. Conclusion
We present the first exploration of structure- and appearance-controllable image stylization. Our contributions include DiffArtist, a styler that fully disentangles structure and appearance during the diffusion process, and a human-aligned evaluator to assess structural and appearance fidelity at the semantic level. Our extensive analysis proves that semantically-rich representations are essential for both structure and appearance style. We demonstrated that our design allows for high style fidelity and controllability, similar to that of a human artist. We believe the objective established in this paper—to stylize in both structure and appearance—offer a roadmap for the next generation of generative art tools to produce artistically meaningful paintings.
7. Acknowledgment
This research was supported by the Hong Kong Research Grants Council (GRF-15229423).
@@ Appendix A Additional Qualitative Results
Figure 7. Tradeoff between Structure and Appearance Style Control. We present the experimental (solid lines) and simulation (dotted lines) trends of structural and appearance variation in the diffusion process. Due to quadratic growth, highly noisy steps are required to achieve strong appearance styles, which are associated with significant structural variation, which can violate semantics. Top: Example stylization results starting from different steps, using the prompt: “Fauvism painting”. When the appearance strength is high ($t=0.8T$), the structure (legs of the horse) is incorrectly modified.
@@
Visualization Additional appearance stylization is in Fig. 10, the source image is in Fig. 11. A grid of different images with different styles is in Fig. 12. Additional structure control is in Fig. 13 and Fig. 14.
Additional qualitative comparisons on stylization. We show an extended comparison with the previous reference-based method in Fig. 15. More qualitative comparisons with existing text-driven image stylization and manipulation methods can be found in Fig. 16 and Fig. 17.
Additional comparisons on fine-grained control. We provide additional comparisons with other control methods in Fig. 18 and Fig. 19. These results demonstrate the advantage of DiffArtist in providing disentangled structural and appearance-level style control. In particular, the Ctrl-X, as an image editing method, produces less visually pleasing results when applied to image stylization. This is because they have a different definition of appearance and structure for editing real photos.
@@ B.1. Theoretical analysis
(7) $\mathbf{x}_{t-1}$ $=\mathcal{A}_{t}\mathbf{x}_{t}+\mathcal{B}_{t}\epsilon_{\theta}(\mathbf{x}_{t},t;y),\quad\mathcal{A}_{t}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}$
$\mathcal{B}_{t}$ $=\sqrt{1-\alpha_{t-1}}-\frac{\sqrt{\alpha_{t-1}(1-\alpha_{t-1})}}{\sqrt{\alpha_{t}}}$
@@
where $\epsilon^{\prime}$ denotes an ideal denoiser that perfectly models the transition distribution $q(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$. In Eq. 8, the conceptual denoising term from $\mathbf{x}_{T}$ to $\mathbf{x}_{\tau}$ preserves the structure and appearance in $I_{c}$. The stylization trajectory from $\mathbf{x}_{\tau}$ to $\mathbf{x}_{t}$ is what is actually computed, which introduces the desired appearance based on the prompt $y$ but may also lead to uncontrolled structure generation. In this paradigm, preserving original appearance and generating new structure are undesirable side effects. Intuitively, a larger $\tau$ leads to a shorter trajectory of $\mathbf{x}_{T}\to\mathbf{x}_{\tau}$, resulting in a stronger appearance in $\hat{\mathbf{x}}_{0}$ with weaker structure preservation (uncontrolled structure stylization), while a lower $\tau$ sacrifices stylization strength for stronger structure preservation. In other words, one can not arbitrarily control the strength of appearance and structure without affecting the other.
By combining Eqn. 2 and Eqn. 8, we can quantitatively assess the strength of content preservation and stylization in the diffusion process under a particular noise schedule $\alpha_{1:T}$. Specifically, we further assume the SNR of structure and appearance has a linear relationship with that of the $\mathbf{x}_{t}$ for $t$:
@@ B.2. Simulation
To derive the theoretical trends of structure and appearance strength during the diffusion process, we introduce an additional assumption. Specifically, we assume that the relative significance of each unweighted denoising step $\epsilon(\mathbf{x}_{t},t;y)$ on the final stylized image remains consistent across different timesteps for both structure and appearance. In other words, we assume the SNR of structure and appearance has a linear relationship with that of the $\mathbf{x}_{t}$ at time $t$ as characterized by the noise schedule $\alpha_{1:T}$:
@@
With the above assumption, the effect of varying $\tau$ on the structure and appearance of the final stylized image could be derived in closed form. In practice, we use the following code to calculate iteratively:
@@ B.3. Empirical Result
Due to the inherent inaccuracy of DDIM inversion, the estimation of $\mathbf{x}_{\tau}$ may be imperfect, resulting in unintended modifications in the final sampled image even if no style prompt $y$ is used. To address this issue, we adopt an alternative strategy by randomly sampling 500 Gaussian noise as the $\mathbf{x}_{T}$ of content, which are paired with 500 content prompts. We treat the images denoised using content prompts for $\tau=T$ steps as the content image, which simulates a perfect inversion technique. To stylize an image, we first denoise the $\mathbf{x}_{T}$ with the content prompt for $\tau$ steps, which is subsequently denoised with the style prompt for $T-\tau$ steps. The LPIPS between the stylized image and the content image is used as the empirical structural strength. In contrast, the CLIP Deception score (correct classification rate among a set of styles) is used as the empirical appearance strength. The following 10 style prompts are used:
@@
The results for both simulation and empirical results are in Fig. 7. The result shows a good fit, and it turns out that the structure modification appears to be linear, with the stylization strength being quadratic with respect to $\tau$. Moreover, this result further evidenced the issue of S-A entanglement in the diffusion process.
@@ C.1. Implementation Details
The stylized images, style prompt, and the instruction prompts are fed to MLLM for inference. We compose stylized images as a grid image with numbers at the top-left corner. The full prompt template for structure and appearance score is available in Tab. 6.
@@ C.2. Correlation with Human Preference
Human Question Collection We distributed the questionnaire on a crowd-sourcing platform, where each participant was required to complete up to 20 randomly sampled ranking tasks. An example of the user interface is provided in Fig. 20. A total of 200 participants took part in this study. To ensure the validity of the responses, we included attention-check questions. If a participant answers an attention-check question incorrectly, all of their responses will be marked invalid. Responses that are made with less than 20 seconds are also removed.
@@ D.2. Additional Ablations on S2A Injection
In this section, we provide additional ablations to study the effect of proposed S2A design.
@@ D.3. Failure Case
In our experiment, we identify a rare (<<1%) and special failure case in the proposed methods. Specifically, for certain content image, its stylization result will consistently contains black and white chessboard-pattern artifacts. We provide one example in Fig. 9.